OpenStudy (anonymous):

tell whether graph opens upward or downward. then find the axis of symmetry and vertex of graph of function y=-x^2+9

OpenStudy (anonymous):

coeff of x^2 = -ve Therefore, it opens downward. axis of symmetry = -b/2a = 0 vertex = (0,9)

OpenStudy (callisto):

let's consider the graph having a form y= a(x-h)^2 +k If a>0, it opens upwards and vice versa Axis of symmetry is x=h vertex is (h,k) consider your case, y=-x^2+9 can be rewritten as y= (-1)(x-0)^2 +9 So a<0, it opens downwards h=0, so axis of symmetry is x=0 k=9, so the vertex is at (0,9)